Two positive point charges are a fixed distance apart. The sum of their charges is QT. What charge must each have in order to (a) maximise the electric force between them and (b) minimise it?



1
Expert's answer
2021-02-07T19:09:12-0500

a)


"F=k\frac{q\cdot(Q_T-q)}{r^2}"


In order for the force to be maximal, the condition "\frac{dF}{dq}=0" must be fulfilled.


"\frac{dF}{dq}=0\to \frac{d}{dq}(qQ_T-q)=0\to q=Q_T/2"


"q_1=q_2=Q_T\/2" . Answer


b)


If "q_1=Q_T" and "q_2=0" then the force between them will be zero, which is the minimum possible force. Answer